find vertex of 3x^2+x-2=0 then graph it
Answer by stanbon(75887)

find vertex of 3x^2+x-2=0 then graph it
---
Vertex occurs where x = -b/(2a) = -1/6
f(-1/6) = 3(1/36)+(-6/36)-(72/36)
------
= (3-6-72)/36 = -75/36
======
Vertex:: (-1/6 , -75/36)
-----------------------
Cheers,
Stan H.
===============